### v1

# Filter Like You Test: Data-Driven Data Filtering for CLIP Pretraining

Abstract
We introduce Filter Like You Test (FLYT), a method for curating large-scale vision-language datasets that learns the usefulness of each data point as a pretraining example. FLYT trains a scoring model that learns to weigh each example using gradient signals from downstream tasks training sets. Using the same training methodology, we develop Mixing-FLYT (M-FLYT), which takes the per-example scores generated by different scoring methods and learns to unify them into a single score. Our training methodology naturally produces a distribution over the training examples, which we leverage through Soft Cap Sampling (SCS), a strategy for obtaining a filtered pretraining dataset from per-example probabilities that samples examples while preventing over-representation through a repetition penalty. Using all three methods, we achieve 40.1% ImageNet zero-shot accuracy on the DataComp medium scale filtering benchmark, a 1.9% absolute accuracy increase over all previous results and a 5.5% increase over results that—like us—use only public resources.
1 Introduction
[…]
Figure 1: In each FLYT training loop, the scoring model takes a batch of upstream data and generates a score for each example. These scores are converted to weights via softmax. A reference model processes the upstream batch, and the resulting embeddings, together with the weights, are used to compute a weighted CLIP loss. The reference model is then updated using gradients from this loss. Next, this updated reference model processes downstream data to compute a downstream loss, which produces a gradient signal that passes through the updated reference model all the way to the scoring model parameters.
Filtering the pretraining data of CLIP models [33] has a dramatic effect on the resulting model’s quality [36, 37, 15, 48]. Despite intense research and considerable progress, the best existing methods select data based on heuristic proxies for usefulness [14, 20, 46, 27, 25, 19]. In this work, we take a different approach: adhering to the core principle of machine learning, we optimize a loss function that measures the usefulness of data for pretraining.
We introduce three methods that progressively build upon each other to achieve state-of-the-art performance on the DataComp medium scale filtering benchmark [15]. At the core is Filter Like You Test [^1] The name is inspired by FLYP [16]. (FLYT), a method for training a scoring model that predicts the weight an image-text pair should have in CLIP pretraining. Mixing-FLYT (M-FLYT) applies the same training process to merge existing scoring methods. The third method, Soft Cap Sampling (SCS), is a simple sampling strategy that leverages the probabilistic structure of FLYT’s scores to construct the final pretraining dataset.
Filter Like You Test (FLYT) (Section 3.2).
FLYT recasts data selection as data re-weighting [34], and learns to upweight (and hence select) the pretraining data that contributes the most to the model’s downstream performance. We jointly train two models: a scoring model that evaluates the usefulness of training examples, and a reference model that provides feedback on these evaluations. During training, the scoring model assigns weights to each example in the batch, which are then used in our weighted CLIP loss function to update the reference model. The updated reference model is then evaluated on downstream data, allowing gradient signal from the downstream loss to flow through it, to the weights, and finally back to the scoring model. This allows the scoring model to give higher weights to examples that reduce the downstream loss of the reference model. We then use the trained scoring model to generate per-example scores. Figure 1 illustrates a single training step of FLYT and Algorithm 1 provides pseudocode. When applying on fixed image/text embeddings, FLYT achieves 31.6% zero-shot ImageNet accuracy, and 32.3% on the average of 38 DataComp evaluation tasks, comparable to the best baseline from Gadre et al. [15].
Mixing-FLYT (M-FLYT) (Section 3.3).
Previous work aggregates different filtering methods to obtain the best results. They either sum together different scores [20], take the intersection/union of the filtered datasets [15, 27, 20, 46] or pipeline the filtering methods one after the other [49, 46]. Departing from heuristic approaches, we leverage FLYT to directly optimize over a space of mixing methods defined by a simple linear combination of scores. Replacing the aggregation method of [20] with M-FLYT improves ImageNet accuracy by 1.3%. More importantly, M-FLYT can aggregate more scores. When used with 12 different scores, it outperforms the best single score by 1.7% on ImageNet accuracy, and by at least 0.8% on average accuracy.
Soft Cap Sampling (SCS) (Section 3.4).
FLYT produces a probability distribution over the training examples, allowing us to obtain a filtered subset by simply sampling from the distribution. However, sampling with replacement produces too many repetition of the top-scoring examples, and sampling without replacement produces too few. To interpolate between the two sampling methods we propose SCS, a strategy that samples with replacement but reduces repetition using a penalty term (see Algorithm 2). Using only publicly available resources, M-FLYT with SCS has 40.1% ImageNet zero-shot accuracy, which is a 1.9% absolute increase compared to the top entry in the DataComp leaderboard, and a 5.5% improvement compared to the best entry that uses only publicly available resources. Our filtered dataset also achieves 37.5% average accuracy across all evaluation tasks, which is within 1.3% of the best in the leaderboard, and a 0.2% improvement over the best entry using only publicly available resources. [^2] This comparison is made against results reported in the DataComp leaderboard, some of used the full 128M DataComp medium candidate pool, while we were only able to download 119M examples. Table 1 compares our filtered dataset to filtered datasets of prior work when constrained to our downloaded data. On the DataComp small scale (which we did not evaluate until the very end of the project) our method sets a new state-of-the-art in both ImageNet and average accuracy (see Table 3).
Input: Scoring model ${q}_{{\phi}_{0}}$ with initial parameters ${\phi}_{0}$, reference model ${f}_{{\theta}_{0}}$ with initial parameters ${\theta}_{0}$, upstream dataset ${D}$ and downstream dataset ${C}$
Parameters: Number of steps $T$, upstream data batch size ${B}$, downstream data batch size ${B^{\prime}}$, weighted loss ${\ell_{\mathrm{WCL}}}$, and downstream loss ${\ell_{\mathrm{DS}}}$
1 for $t=0$ to $T-1$ do
2       ${{{z}_{1:{B}}}}\sim{D}$, ${{\hat{z}}_{1:{B^{\prime}}}}\sim{C}$
3       ${{s}_{1:{B}}}={q}_{{\phi}_{t}}({{z}_{1:{B}}})$
4       ${{w}_{1:{B}}}=\mathrm{softmax}({{s}_{1:{B}}})$
5       ${L_{\mathrm{up}}}={\ell_{\mathrm{WCL}}}({f}_{{\theta}_{t}}({{z}_{1:{B}}}),{{w}_{1:{B}}})$
6       ${\theta}_{t+1}={\mathrm{update}}(\nabla_{{\theta}_{t}}{L_{\mathrm{up}}})$
7       ${\hat{L}_{\mathrm{down}}}={\ell_{\mathrm{DS}}}({f}_{{\theta}_{t+1}}({{\hat{z}}_{1:{B^{\prime}}}}))$
       ${\phi}_{t+1}={\mathrm{update}}(\nabla_{{\phi}_{t}}{\hat{L}_{\mathrm{down}}})$
        $\triangleright$ see Appendix A
8
return ${\phi}_{T}$
Algorithm 1 Filter Like You Test (FLYT)
Input: Training examples ${z}_{1:{M}}$ and their scores ${s}_{1:{M}}$
Parameters: Repetition penalty ${\alpha}$, target dataset size ${N}$ and sample count per iteration ${G}$.
1 ${\hat{D}}=\left[\text{ }\right]$ while $\lvert{\hat{D}}\rvert<{N}$ do
2       ${w}_{1:{M}}=\mathrm{softmax}({s}_{1:{M}})$
3       $i_{1:{G}}=$ Sample ${G}$ indices according to distribution ${w}_{1:{M}}$ without repetitions.
4       ${\hat{D}}=\mathrm{concatenate}({\hat{D}},{z}_{i_{1:{G}}})$
5       ${s}_{i_{1:{G}}}={s}_{i_{1:{G}}}-{\alpha}$
return ${\hat{D}}$
Algorithm 2 Soft Cap Sampling (SCS)
Data and code release.
We share all the code required to reproduce our work, including FLYT, M-FLYT training, and SCS sampling at https://github.com/formll/FLYT. Additionally, we publish our best-performing scoring and mixing models along with the resulting filtered datasets, and the M-FLYT input scores at https://huggingface.co/collections/formll/flyt-67bb167366ec0fa0d5b8e4bd.
2 Related work
Early work on curating large-scale image-text datasets established methods to extract, filter, and organize pairs of images and captions from the web [38, 4, 42, 12]. Beyond assembling large collections, several lines of research address dataset pruning and deduplication [41, 1], example characterization [26, 44], and quality assessment [32].
CLIP [33] leveraged large image-text datasets via contrastive learning, producing the first generation of vision-language foundation models and revolutionizing computer vision. While the original CLIP was trained on a proprietary set of 400M image-text pairs, subsequent efforts introduced open-source alternatives. LAION-400M [36] and LAION-5B [37] filtered large web-crawled corpora primarily using the cosine similarity between image and text embeddings from a pretrained CLIP model, along with basic image and text quality metrics. Xu et al. [48] attempted to reproduce CLIP’s original data pipeline and managed to outperform the original model.
Recognizing the importance of data quality for CLIP training, Gadre et al. [15] introduces the DataComp benchmark. DataComp fixes the data pool as well as the training and evaluation procedures, and allows researchers to focus on comparing filtering methods while keeping all other parameters fixed. Like us, many papers use the DataComp benchmark to evaluate their filtering methods [49, 27, 20, 14, 46, 25]. Below, we survey the literature on data filtering for DataComp.
Filtering methods.
Gadre et al. [15] consider several baseline filtering methods, and obtain their best results with a combination of CLIP score and a clustering-based filtering method using similarity to the ImageNet training set. T-MARS [27] observe that images containing texts that overlap with the caption received high CLIP scores, while not contributing much to the model training. To address this issue, they mask out these texts before calculating CLIP score. For better results, they take the intersection of this approach and their variation of Maini et al. [26]. Yu et al. [49] develop a sequential pipeline that combines many filtering techniques, which they divide to three approaches: single modality filtering, cross modality filtering, and distribution alignment. DFN [14] use CLIP score from a model fine-tuned on a large high-quality private data. HYPE [20] suggest that relying only on the image-text alignment information produced by CLIP score is insufficient, and propose to also consider the individual information of the images and text. They use hyperbolic embeddings to generate a hyperbolic similarity score in addition to image and text specificity scores, which they sum with the traditional Euclidean CLIP score and the image-based score. NegCLIPLoss [46] introduce a variation to the standard CLIP score that normalizes it by the similarity score of the example with other examples in a batch, and after filtering using this score, they filter the remaining dataset using a downstream examples similarity metric. While these heuristic approaches are successful, our approach aims to directly learn how to estimate an example’s quality.
Data sampling.
Given a scoring of the data pool, most filtering approaches select the top 15-20% of examples, resulting in each example appearing 5–7 during training in the DataComp benchmark. Goyal et al. [17] analyzed the quality-quantity trade-off of dataset filtering, and found that for the compute budget of the medium scale DataComp benchmark selecting the top 20% of examples is optimal, with more aggressive filtering leading to diminishing returns. However, their analysis only considers uniform sampling while top performing methods created non-uniformly sampled datasets through concatenation of multiple subsets of data. In their work, HYPE [20] concatenate their selected subset with DFN’s subset, while NegCLIPLoss [46] concatenate their subset with both HYPE and DFN. This implicitly increases the sampling frequency of examples selected by multiple methods, with examples chosen by all three methods being sampled approximately 10 times during training. Our proposed SCS method extends this approach of non-uniform sampling using a simple probabilistic procedure; in our top performing dataset, individual examples appear up to 56 times.
Learned data weighting.
Ren et al. [34] introduced an algorithm for online example weighting which utilizes gradient signals from a validation dataset for producing per-example weights. Shu et al. [40] employs a similar idea to Ren et al. [34] with the addition of a weighting model that is optimized to weigh the examples. FLYT is similar to these methods in that it too learns weights and backpropogates through a “target loss” after a gradient update on the “training loss.” However, unlike FLYT, these methods consider online data weighting for standard supervised learning, with the intention of improving robustness to noise and data imbalance, mostly in small-scale settings.
3 Method
In this section, we describe FLYT in detail. We introduce our notation by going over standard CLIP training (Section 3.1), then move to describing FLYT (Section 3.2), M-FLYT (Section 3.3), and SCS (Section 3.4). Finally, we cover the weighted CLIP loss we use to train FLYT and M-FLYT (Section 3.5).
3.1 Notation and problem setting
CLIP training.
We denote a CLIP training example ${z}=({{z}^{I}},{{z}^{T}})\in{D}$ where ${D}$ is our pretraining dataset and ${{z}^{I}},{{z}^{T}}$ are the training image and text, and denote a batch of ${B}$ training examples by ${{z}_{1:{B}}}$. We train a CLIP model with parameters ${\theta}$ that maps each example ${z}$ to features ${{f}_{{\theta}}({z})}=({{f}_{{\theta}}({{z}^{I}})},{{f}_{{\theta}}({{z}^{T}})})$. Letting ${{f}_{{\theta}}({{z}_{1:{B}}})}$ denote the features of all the examples in the batch, the standard CLIP loss is
${\ell_{\text{C}}}\left({{f}_{{\theta}}({{z}_{1:{B}}})}\right)=\frac{{\ell_{\text{image}}}+{\ell_{\text{text}}}}{2},$ (1)
where
${\ell_{\text{image}}}$ $={\ell_{\text{contrast}}}({{f}_{{\theta}}({{z}^{I}})},{{f}_{{\theta}}({{z}^{T}})}),$
${\ell_{\text{text}}}$ $={\ell_{\text{contrast}}}({{f}_{{\theta}}({{z}^{T}})},{{f}_{{\theta}}({{z}^{I}})}),$
${\ell_{\text{contrast}}}(u_{1:B},v_{1:B})$ $=\sum_{i=1}^{{B}}-\log\left(\frac{e^{{\tau}\langle u_{i},v_{i}\rangle}}{\sum_{j=1}^{{B}}e^{{\tau}\langle u_{i},v_{j}\rangle}}\right),$
and ${\tau}$ is the learnable temperature parameter.
3.2 Filter Like You Test (FLYT)
In this section we describe FLYT. Our goal is to learn a scoring model ${q}_{\phi}:{D}\to\mathbb{R}$, parameterized by ${\phi}$, that maps each example ${z}\in{D}$ to a score ${s}$, such that the higher ${s}$ is, the more useful ${z}$ is as a pretraining example. We overload notation to let
${{s}_{1:{B}}}={{q}_{\phi}}({{z}_{1:{B}}})$
denote the scores computed element wise over the batch. We obtain weights over the batch ${{w}_{1:{B}}}$ using a softmax transformation:
${\mathrm{softmax}}({s}_{i})=\frac{\exp({{s}_{i}})}{\sum_{j=1}^{{B}}\exp({{s}_{j}})},$
${{w}_{1:{B}}}={\mathrm{softmax}}({{s}_{1:{B}}}).$
Now consider a reference CLIP model ${f}_{{\theta}}$. We define a weighted CLIP loss
${L_{\mathrm{up}}}={\ell_{\mathrm{WCL}}}({{f}_{{\theta}}({{z}_{1:{B}}})},{{w}_{1:{B}}})$
which is a function of both the reference model embeddings, and the weights produced by the scoring model. This loss serves as an approximation to the loss we would have gotten had we used the scoring model to filter the batch. We discuss the weighted CLIP loss in detail in section Section 3.5.
We can use this loss and its gradient to update ${\theta}$; let
${{\theta}_{+}}({{w}_{1:{B}}})={\mathrm{update}}(\nabla_{{\theta}}{L_{\mathrm{up}}})$
be the result of a gradient-based update to ${\theta}$. For example, using SGD with learning rate $\eta$, we have ${{\theta}_{+}}({{w}_{1:{B}}})={\theta}-\eta\nabla_{{\theta}}{L_{\mathrm{up}}}$. Using a more sophisticated optimizer like AdamW, ${{\theta}_{+}}({{w}_{1:{B}}})$ takes a more complicated form, but we can still think about it as a function of ${{w}_{1:{B}}}$.
Next, we introduce the key part of our approach, a batch of downstream data (e.g., the ImageNet training set). Let ${{\hat{z}}_{1:{B^{\prime}}}}$ denote this batch (note that ${B^{\prime}}$ does not necessarily have to be equal to ${B}$). We use this batch to obtain a gradient signal for updating ${{w}_{1:{B}}}$ and hence the scoring model parameters ${\phi}$. To that end, define:
${\hat{L}_{\mathrm{down}}}={\ell_{\mathrm{DS}}}({f}_{{{\theta}_{+}}({{w}_{1:{B}}})}({{\hat{z}}_{1:{B^{\prime}}}})),$ (2)
Where ${\ell_{\mathrm{DS}}}$ is some loss over the downstream data: we experiment with standard CLIP loss ${\ell_{\text{C}}}$, as well as cross-entropy (CE), with or without temperature; see Section C.1 for details. We can then update ${\phi}$ using $\nabla_{\phi}{\hat{L}_{\mathrm{down}}}$.
Finally, we generate per-example scores using our trained scoring model and use these scores for data filtering.
3.3 Mixing-FLYT (M-FLYT)
Inspired by past success in mixing multiple filtering heuristics [15, 20, 46, 49], we consider another application of FLYT.
Instead of training a scoring model ${q}:{D}\to\mathbb{R}$ that predicts the per-example scores directly from ${z}$, we take advantage of other scoring methods and train what we call a mixing model. Our mixing model ${p}:\mathbb{R}^{{k}}\to\mathbb{R}$ takes as input ${k}$ different scores, generated by ${k}$ different scoring methods, and outputs a single unified score ${s}$. Other than the different input of the mixing model, we train M-FLYT in the exact same way as FLYT.
Using M-FLYT gives us the advantage of not needing to learn example importance from scratch, and instead we use the information gained by other scoring methods. In fact, we notice that a simple linear mixing model is enough for obtaining strong results.
3.4 Soft Cap Sampling (SCS)
We now consider the task of obtaining a filtered pretraining dataset from per-example scores. A principled data filtering approach should use these scores for deciding the number of repetitions of each example in the filtered dataset. It is not clear how to do that when using a similarity based score as in most prior work. In contrast, a sampling strategy emerges directly from the definition of FLYT, since the scores are naturally interpreted as the learned (log) probability of including an example in the batch.
A straightforward approach would be to sample from this distribution 128M times to create our final dataset. In practice, doing that results in over-represented examples. To correct this, we introduce SCS (Soft Cap Sampling): every time we add an example to our filtered data, we subtract a “repetition penalty” $\alpha$ from its score, reducing the probability it is sampled again. For implementation efficiency, we perform the sampling procedure iteratively. In each iteration, we sample a batch of ${G}=100K$ examples without repetitions, then subtract ${\alpha}$ from the score of each batch member. Algorithm 2 presents SCS, and Figure 2 illustrates the distribution generated via SCS.
Figure 2: Histogram of example repetitions using SCS on the probabilities produced by M-FLYT. See Section 4.4 for more details.
3.5 Weighted CLIP loss
Our definition of weighted CLIP loss follows these guiding principles:
  1. 1.
    Most importantly, a larger weight means the corresponding example is more represented in the loss.
  2. 2.
    Setting an example’s weight to 0 should be equivalent to excluding it from the batch.
  3. 3.
    Uniform weights should result in the standard CLIP loss.
  4. 4.
    The weight should reflect the example’s importance both as a positive and as a negative example.
To realize these principle we modify the standard CLIP loss Equation 1, showing the modifications in red:
${\ell_{\mathrm{WCL}}}\left({{f}_{{\theta}}({{z}_{1:{B}}})},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{w}_{1:{B}}}}\right)=\frac{{\ell^{\mathrm{w}}_{\text{image}}}+{\ell^{\mathrm{w}}_{\text{text}}}}{2},$ (3)
where
${\ell^{\mathrm{w}}_{\text{image}}}={\ell^{\mathrm{w}}_{\text{contrast}}}\left({{f}_{{\theta}}({{z}^{I}})},{{f}_{{\theta}}({{z}^{T}})},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{w}_{1:{B}}}}\right),$
${\ell^{\mathrm{w}}_{\text{text}}}={\ell^{\mathrm{w}}_{\text{contrast}}}\left({{f}_{{\theta}}({{z}^{T}})},{{f}_{{\theta}}({{z}^{I}})},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{w}_{1:{B}}}}\right),\mbox{and}$
${\ell^{\mathrm{w}}_{\text{contrast}}}\left(u_{1:B},v_{1:B},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{w}_{1:{B}}}}\right)$
$\qquad=\sum_{i=1}^{{B}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}}\log\left(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}}e^{{\tau}\langle u_{i},v_{i}\rangle}}{\sum_{j=1}^{{B}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{j}}e^{{\tau}\langle u_{i},v_{j}\rangle}}\right).$
4 Experiments
Table 1: Comparison of our method to top performing methods on the DataComp medium scale filtering benchmark. DFN-FT is our best reproduction of DFN [14]. The mark * denotes results from the original paper, and $+$ indicates evaluations using our downloaded 119M candidate pool using the authors’ provided UIDs. Table 13 compares our evaluations to the results reported in the corresponding papers.
Filtering Method Public Resources ImageNet ImageNet dist. shifts VTAB Retrieval Average
[…]
DFN-FT [our reproduction of 14] Yes 0.342 0.274 0.357 0.290 0.348
HYPE 20% [20] Yes 0.338 0.269 0.357 0.286 0.343
NegCLIPLoss+ [46] Yes 0.333 0.273 0.361 0.251 0.352
DFN+ [14] No 0.367 0.303 0.371 0.280 0.364
HYPE+DFN+ [20] No 0.372 0.304 0.374 0.287 0.368
DFN+NegCLIPLoss+ [46] No 0.371 0.304 0.390 0.285 0.378
HYPE+DFN+NegCLIPLoss+ [46] No 0.381 0.310 0.392 0.284 0.378
FLYT Yes 0.316 0.251 0.331 0.276 0.323
M-FLYT Yes 0.359 0.294 0.383 0.310 0.371
M-FLYT+SCS (${\alpha}=0.15$) Yes 0.401 0.311 0.391 0.292 0.375
M-FLYT+SCS (${\alpha}=0.25$) Yes 0.395 0.309 0.393 0.310 0.378
In this section, we present the steps we took to create our top performing dataset which achieves 40.1% zero-shot ImageNet accuracy on the DataComp medium scale benchmark. We first review key details about our training and evaluation setup (Section 4.1), then discuss the challenges, design choices and results for FLYT (Section 4.2), M-FLYT (Section 4.3), and SCS (Section 4.4). Table 1 summarizes our findings and compares them to the state-of-the-art.
4.1 General Setup
We experiment mainly with the medium scale DataComp filtering track benchmark [15]. In this benchmark, the goal is to filter a 128M example candidate pool sourced from Common Crawl [^3] https://commoncrawl.org. The filtered data is then used to train a ViT-B/32 CLIP using a fixed procedure, and the resulting model undergoes a fixed evaluation on 38 downstream tasks, with performance typically reported in four main categories: ImageNet, ImageNet distribution shifts, VTAB, and retrieval, along with the overall average across all tasks.
For our experiments we were only able to download 119M out of the 128M DataComp candidate pool. To compare our filtered datasets to those published in prior work, we reapply the DataComp training and evaluation procedure on the intersection of the published datasets and our downloaded pools. As shown in Table 13, the accuracy degradation due to the missing data is between 0.1% and 1%.
The fixed DataComp medium training procedure goes through 128M input examples (hence using a fixed amount of compute) regardless of the size of the filtered data; if the filtered dataset has less than 128M examples, it is reshuffled and repeated until reaching 128M examples. In all experiments except those specifically evaluating sampling strategies, we follow the standard strategy of selecting the top 20% scoring examples, which means each example is seen by the model 5 or 6 times during training. Using SCS, we control how many times each example will be seen, forming a dataset with repeating data and using a shuffle buffer to spread data repetitions across training.
For both FLYT and M-FLYT, we use a CLIP ViT-B/32 model as the reference model. Unless otherwise mentioned, we use the unfiltered DataComp medium pool as the upstream dataset, and the ImageNet training set as the downstream dataset.
To enable large-scale training of FLYT, we develope a data parallelism approach, which is detailed in Appendix A.
4.2 FLYT Results
Our scoring model consists of CLIP ViT-B/32 [33] base and a ${\mathrm{FFN}_{\mathrm{GLU}}}$ [39] head. We initialize both the scoring model base and the reference model with OpenAI’s checkpoint and keep the base frozen when training. We outline key decisions here, and give detailed training parameters and ablations in Appendix C.
We evaluate three alternatives for the downstream loss function: The CLIP loss, as defined in Equation 1 ${\ell_{\text{C}}}({f}_{{{\theta}_{+}}({{w}_{1:{B}}})}({{\hat{z}}_{1:{B^{\prime}}}}))$, standard cross entropy (CE) loss, and CE with a temperature parameter (see definitions in Section C.1). Our experiments show that CLIP loss and CE with temperature demonstrate similar performance, both outperforming the standard CE loss (Table 8). For subsequent experiments, as well as the FLYT result reported in Table 1, we use CE with temperature because it better resembles the accuracy metric used to evaluate most DataComp tasks. We found it essential to use separate learned temperature parameter for the upstream and downstream loss functions when using either the CLIP or CE with temperature loss. We initialize the downstream temperature parameter ${\tau_{\mathrm{DS}}}$ to 1/0.07, as is the default when training a CLIP model from scratch, while the initial value for the upstream temperature is set to 100 as part of the pretrained reference model weights. The downstream temperature is updated during training using $\nabla_{{\tau_{\mathrm{DS}}}}{\hat{L}_{\mathrm{down}}}$.
We tried several alternatives to our base configuration. Training the full scoring model including the ViT-B/32 base lowered the performance significantly, as did freezing the reference model during training. Using weaker models to initialize either the reference model or scoring model base produced worse results. Initializing the reference model with a stronger model has negligible effect, while initializing the scoring model base with a stronger model yielded improved results (Table 8). We also tried expanding the downstream data beyond ImageNet by including training sets from 22 DataComp tasks, but observed no significant difference in performance (Table 9). As discussed in Section 5, we believe that our FLYT training stack has significant room for improvement.
4.3 M-FLYT Results
Table 2: Performance and linear weights of each input to M-FLYT, along with baseline experiments. For HYPE [20] and NegCLIPLoss [46], we replicate the scores and mixing methodology used in the original papers.
Input Linear weight ImageNet Average
ViT-B/32 0.08 0.282 0.327
ViT-L/14 0.21 0.267 0.317
DFN-Base 0.60 0.297 0.333
DFN-FT 0.80 0.342 0.348
HYPE ${-d_{L}}$ 0.45
HYPE ${\epsilon_{i}}$0.05
HYPE ${\epsilon_{t}}$ 0.02 0.309 0.341
NegCLIPLoss 0.51
[…]
IN1K-Classifier 0.36 0.268 0.294
CC2M-Classifier 0.17 0.287 0.308
FLYT 0.65 0.316 0.323
M-FLYT 0.359 0.371
Sum 0.324 0.339
Standardized Sum 0.344 0.361
For M-FLYT, we train a linear mixing model using input scores that are standardized to zero mean and unit standard deviation. The standardization both improves performance and provides interpretable weights for each input method. For M-FLYT we empirically found that initializing the reference model with a CLIP model trained using DataComp’s medium scale configuration on their unfiltered dataset and using a downstream CE loss without temperature produced better results: see Section C.6
[…]
We compare M-FLYT to its individual input components and to two baseline approaches. The first baseline uses the sum of all input scores, while the second sums the scores after standardization to zero mean and unit standard deviation. Table 2 demonstrate that M-FLYT outperforms all individual scores as well as both baseline aggregations. Notably, while FLYT has modest performance as a standalone method, it receives the second-highest weight, surpassed only by DFN-FT, indicating that FLYT picks up scoring signal that are complementary to those found in prior work. Table 11 demonstrates that excluding any individual input score from M-FLYT’s training leads to performance degradation, suggesting that the method effectively leverages multiple signals, and could potentially achieve stronger performance given more (or better) input scores.
4.4 SCS Results
Figure 3: Comparison of sampling strategies (SCS, HCS, standard threshold filtering, and No Cap) showing their effect on ImageNet accuracy (top) and average accuracy (bottom). SCS was tested with ${\alpha}$ values from 0.1 to 0.6, HCS with ${\beta}$ values from 5 to 25 and standard threshold with values from 5% to 25%
For SCS, we always sample in batches of ${G}=100K$, and experiment with a repetition penalty ${\alpha}$ between 0.1 and 0.6. When preparing the filtered dataset, we use a shuffle buffer of size 1.5M in the DataComp resharder to prevents batches from containing many copies of the same example.
When using the scores produced by M-FLYT, SCS outperforms every other filtering method in the DataComp ImageNet Leaderboard (Table 1). Figure 3 shows the influence of the repetition penalty parameter ${\alpha}$, and compares SCS to standard threshold filtering, regular sampling (No Cap), as well as to a sampling method we call Hard Cap Sampling (HCS). HCS allows an example to be sampled until it reaches a “frequency limit” ${\beta}$, and then its sampling probability is set to 0. We note that while SCS outperforms HCS, both methods outperform the standard threshold filtering.
When using a repetition penalty ${\alpha}$ of 0.15, which yields 25.67M unique examples, we achieve the highest ImageNet accuracy. However, increasing ${\alpha}$ to 0.25 generates more unique examples (29.43M) and increases average accuracy across all 38 downstream tasks. This pattern, where example diversity improves average accuracy but reduces ImageNet accuracy, also appears in threshold filtering: selecting the top 10% of examples (11.93M) performs better on ImageNet compared to selecting the top 20% (23.86M), but results in lower average accuracy across tasks.
4.5 Small Scale Results
Table 3: Comparing our methods to top performing methods on the DataComp small scale filtering benchmark. * denotes results we got from the DataComp leaderboard that did not provide a paper to cite.
Filtering Method ImageNet Average
EcoDatum 0.053 0.182
WS [19] 0.056 0.180
HYPE [20] 0.063 0.180
CAM $\cap$ Flipped CLIP 0.064 0.178
M-FLYT 0.072 0.185
M-FLYT + SCS (${\alpha}=0.5$) 0.080 0.185
We evaluate our method on the small scale DataComp benchmark. We use the mixing model trained by M-FLYT described in Section 4.3, and sample using SCS with a repetition penalty ${\alpha}=0.5$ and a batch size ${G}=10K$. Table 3 shows that our method achieves 8% zero-shot ImageNet zero-shot accuracy and 18.5% average accuracy, outperforming the top methods in the DataComp leaderboard on both categories. These results suggest that M-FLYT and SCS can be effectively used across different computational scales. Figure 4 shows the effect of different ${\alpha}$ values
5 Limitations and Future Work
Computational scale is a clear limitation of our work: we only had the resources to experiment at the DataComp medium scale, and could not test our methods on the larger scales. However, in prior work progress on medium scale usually translated to progress on larger scales, which makes us optimistic that our proposed methods will also scale well.
Another notable limitation is that FLYT training on raw inputs did not succeed, and on fixed embeddings did not surpass state-of-the-art methods (though it did contribute to M-FLYT). We believe this is likely a deficiency in our optimization stack, solvable by more careful tuning of learning rate (schedules) for individual components and regularization techniques such as dropout. In similar vein, we believe that more careful parameter tuning should allow us to leverage downstream data that is more diverse than ImageNet.
A third limitation—and direction for future research—is that we use data weighting as a proxy for data selection when defining the FLYT objective function. It would be interesting to try to model data selection more directly via reinforcement learning: at each reference model training step, the scoring model picks a batch of training data and observes a reward signal for the chosen data.
A fourth direction for future work is closing the data selection / pretraining loop by iteratively re-training the scoring model to select the next data to pretrain our model on. FLYT is particularly well-suited to this form of active learning, as it allows us to use the pretraining model as the reference model at each scoring iteration, ensuring that each batch we select is most useful for the model at its current state.
Finally, FLYT should be useful past DataComp and vision-language data filtering. In particular, it will be exciting to extend it to data filtering and mixing for language models.
Acknowledgments
We thank Sonia Lichtenzveig Sela and Alexander Tsverin from the TAU CS systems team for crucial technical support. We thank Alex Fang and Vaishaal Shankar for helpful discussion of the DFN implementation details, and Sanghyuk Chun for sharing the HYPE codebase.
This research was partially supported by the computing resources of the Tel Aviv University Center for AI and Data Science, and the Israeli Science Foundation (ISF) grant no. 2486/21.
Appendix A Implementing data-parallel distribution
CLIP training benefits from a large batch size [33, 6]. FLYT trains a CLIP model while weighting each example compared to others in the batch, which also requires a sufficiently large batch size. However, training FLYT requires more memory than training a standard CLIP model since we have to keep the computational graph from ${\phi}$ to ${\hat{L}_{\mathrm{down}}}$ in order to calculate $\nabla_{\phi}{\hat{L}_{\mathrm{down}}}$. This means that in order to train FLYT on a large scale, we need to implement data-parallel distributed training.
Our goal is to compute $\nabla_{\phi}{\hat{L}_{\mathrm{down}}}$ using data parallelism across multiple GPUs, which is not trivial considering FLYT introduces different dependencies between examples in the batch. We start by using the chain rule to write:
$\nabla_{\phi}{\hat{L}_{\mathrm{down}}}=\left[\frac{\partial{{w}_{1:{B}}}}{\partial{\phi}}\right]\nabla_{{{w}_{1:{B}}}}{\hat{L}_{\mathrm{down}}},$
and focus the second factor
$\nabla_{{{w}_{1:{B}}}}{\hat{L}_{\mathrm{down}}}=\nabla_{{{w}_{1:{B}}}}{\ell_{\mathrm{DS}}}({f}_{{{\theta}_{+}}({{w}_{1:{B}}})}({{\hat{z}}_{1:{B^{\prime}}}})),$
where
${{\theta}_{+}}({{w}_{1:{B}}})={\mathrm{update}}({g}({{w}_{1:{B}}}))$
with
${g}({{w}_{1:{B}}})=\nabla_{{\theta}}{\ell_{\mathrm{WCL}}}({{f}_{{\theta}}({{z}_{1:{B}}})},{{w}_{1:{B}}})$
and ${\mathrm{update}}(\cdot)$ is an optimizer update function. We then use the chain rule again to write:
$\nabla_{{{w}_{1:{B}}}}{\hat{L}_{\mathrm{down}}}=\left[\frac{\partial{g}}{\partial{{w}_{1:{B}}}}\right]\nabla_{{g}}{\ell_{\mathrm{DS}}}(f_{{\mathrm{update}}({g})}({{\hat{z}}_{1:{B^{\prime}}}})),$
We may further apply the chain rule (and the fact the partial derivatives commute) to write
${g}=\sum_{i\in[{B}]}\frac{\partial{\ell_{\mathrm{WCL}}}({f}_{1:{B}},{{w}_{1:{B}}})}{\partial{f}_{i}}\nabla_{{\theta}}{f}_{\theta}({z}_{i})$
where ${f}_{i}$ is shorthand for the embedding ${f}_{{\theta}}({z}_{i})$ that implicitly says we only need to store the value of the embedding and not the computational graph behind it. This way each GPU calculates ${g}$ for its own sub-batch, and we can then gather and sum their result to get ${g}$ for the full batch using the CLIP gradient accumulation technique from Cherti, Mehdi et al. [6]. Using this, we define
${v}\triangleq\nabla_{{g}}{\ell_{\mathrm{DS}}}({f}_{{\mathrm{update}}({g})}({{\hat{z}}_{1:{B^{\prime}}}})),$
which we can again compute using gradient accumulation. Treating ${v}$ as a constant vector independent of ${{w}_{1:{B}}}$, the gradient we would like to calculate is
$\nabla_{{{w}_{1:{B}}}}{\hat{L}_{\mathrm{down}}}=\frac{\partial\langle{g},{v}\rangle}{\partial{{w}_{1:{B}}}}=\frac{\partial}{\partial{{w}_{1:{B}}}}\sum_{i\in[{B}]}\frac{\partial{\ell_{\mathrm{WCL}}}({f}_{1:{B}},{{w}_{1:{B}}})}{\partial{f}_{i}}\langle\nabla_{{\theta}}{f}_{{\theta}}({z}_{i}),{v}\rangle.$
If we define the function
${h}_{i}({\eta})=\frac{\partial{\ell_{\mathrm{WCL}}}({f}_{1:{B}},{{w}_{1:{B}}})}{\partial{f}_{i}}{f}_{{\theta}+{\eta}{v}}({z}_{i})$
Then our gradient is simply
$\nabla_{{{w}_{1:{B}}}}{\hat{L}_{\mathrm{down}}}=\sum_{i\in[{B}]}\frac{\partial}{\partial{{w}_{1:{B}}}}{h}^{\prime}_{i}(0).$
Which is amenable to gradient accumulation.
The first factor
$\frac{\partial}{\partial{\phi}}{{w}_{1:{B}}}=\frac{\partial}{\partial{\phi}}{\mathrm{softmax}}({{q}_{\phi}}({{z}_{1:{B}}}))$
can be calculated separately using gradient accumulation, and we get
$\nabla_{\phi}{\hat{L}_{\mathrm{down}}}=\left[\frac{\partial{{w}_{1:{B}}}}{\partial{\phi}}\right]\nabla_{{{w}_{1:{B}}}}{\hat{L}_{\mathrm{down}}}.$
Appendix B M-FLYT Input Details
In order to experiment with M-FLYT, we reproduce some existing scoring methods and create some new ones:
  1. 1.
    CLIP scores. We use the CLIP scores calculated using OpenAI ViT-B/32 and ViT-L/14 CLIP models [33].
  2. 2.
    DFN. The results of Fang et al. [14] are proprietary and we do not have access to their best DFN model or the scores created by it. The authors provided a version that uses only publicly available resources, but it was not trained using all the improved techniques they discovered. Thus, we reproduced their work using only publicly available resources. We use the CLIP score of 2 DFN networks: one before fine-tuning on ImageNet, and one after. Section B.1 provides the details of our reproduction.
  3. 3.
    HYPE. Kim et al. [20] provided a PyTorch implementation which includes their hyperbolic CLIP weights, and the functionality required to reproduce the image specificity ${\epsilon_{i}}$, text specificity ${\epsilon_{t}}$, and negative Lorentzian distance ${-d_{L}}$. We used these three values as inputs.
  4. 4.
[…]
  5. 5.
    Binary classifiers. We trained two linear binary models that classify between a “high quality” dataset and the DataComp medium scale dataset. We describe them in Section B.2 below.
  6. 6.
    FLYT. We use our best performing FLYT scoring model, as detailed in Section 4.2.
B.1 DFN Reproduction
Table 4: DFN-Base reproduction experiments.
Initialization Num. samples Notes Training dataset ImageNet Average
Random 10M CC12M 0.230 0.300
Random 50M CC12M 0.274 0.324
Random 128M CC12M 0.281 0.328
OpenAI 128M DFN-Base CC12M 0.297 0.333
OpenAI 128M CC12M+CC3M+SS15M 0.286 0.334
OpenAI 512M 2x Batch Size (8192) CC12M 0.296 0.338
Table 5: DFN-FT reproduction experiments
Initialization Num. samples Notes Training dataset ImageNet Average
DFN-Base 400K DFN-FT IN1K+MS COCO+Flickr30k 0.342 0.348
DFN-Base 1.6M IN1K+MS COCO+Flickr30k 0.340 0.348
DFN-Base 3M IN1K+MS COCO+Flickr30k 0.335 0.338
DFN-Base 400K 1/10 LR (5e-5) IN1K+MS COCO+Flickr30k 0.321 0.344
DFN-Base 1M 1/10 LR (5e-5) IN1K+MS COCO+Flickr30k 0.324 0.342
DFN-Base 12M 1/10 LR (5e-5) IN1K+MS COCO+Flickr30k 0.319 0.340
DFN-Base 800K IN1K 0.333 0.343
Since M-FLYT requires good inputs to perform well, we made effort to train a high performing DFN. In the paper, Fang et al. [14] trained a ViT-B/32 CLIP model initialized with OpenAI’s checkpoint on the high-quality private dataset HQITP-350M. Since we do not have access to this dataset, we trained our DFN on CC12M [38]. Similar to Fang et al. [14], we then fine-tune on the combined training sets of IN1K with sampled OpenAI templates as texts, MS COCO, and Flickr30k. Unless otherwise specified, we use the same hyper-parameters as DataComp medium scale training. Tables 4 and 5 show results we obtained with different hyperparameter choices when our reproductions DFN-Base and DFN-FT, respectively.
B.2 Binary Classifiers Input
When training our top performing M-FLYT, we use two binary linear classifiers. These models take as input embeddings generated by OpenAI’s ViT-B/32 model and predict which of two datasets these embeddings come from. IN1K-Classifier classifies between the ImageNet training set and the DataComp medium scale dataset image embeddings. CC2M-Classifier classifies between concatenated image and text embeddings of the DataComp medium scale dataset and CC2M. CC2M is a dataset of 2M examples we created by running our IN1K-Classifier on the CC12M dataset [38] and taking the top 2M examples that the model classified as ImageNet examples. As shown in Table 2, while IN1K-Classifier underperforms compared to CLIP score filtering, it still demonstrates improvement over no filtering and receives substantial weighting from M-FLYT. In contrast, CC2M-Classifier achieves performance similar to CLIP score filtering but receives a lower M-FLYT weighting. We trained them for 40K steps with a batch size of 256, a cosine learning rate with a maximal value of 0.01 and 500 warm-up steps.
Appendix C Experimental Setup
C.1 Downstream loss
We now explain how construct the downstream loss from a downstream batch ${{\hat{z}}_{1:{B^{\prime}}}}$ containing labeled data of the form $\hat{z}_{i}=(x_{i},c_{i})$, where $x_{i}$ is an image and $c_{i}\in\{1,\ldots,{K}\}$ is a class label. First, we uniformly sample templates for each of the classes following DataComp’s class-specific template definitions $t_{1:{K}}=\mathrm{sample}({1:{K}})$. Then, the CE loss is given by
${\ell_{\mathrm{DS}}}({{\hat{z}}_{1:{B^{\prime}}}})=\sum_{i=1}^{{B^{\prime}}}-\log\left(\frac{e^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\tau_{\mathrm{DS}}}}\langle{f}_{{\theta}}(x_{i}),{f}_{{\theta}}(t_{c_{i}})\rangle}}{\sum_{j=1}^{{K}}e^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\tau_{\mathrm{DS}}}}\langle{f}_{{\theta}}(x_{i}),{f}_{{\theta}}(t_{j})\rangle}}\right),$
where ${\tau_{\mathrm{DS}}}$ is a learnable temperature parameter we only use in the CE+temperature variation for the downstream data. Using the standard CE loss, ${\tau_{\mathrm{DS}}}$ is set to 1. Using this notation, we define the downstream CLIP loss mentioned in Section 4.2 with a separate temperature parameter ${\tau_{\mathrm{DS}}}$ as
${\ell_{\text{C}}}({{\hat{z}}_{1:{B^{\prime}}}})=\sum_{i=1}^{{B^{\prime}}}-\log\left(\frac{e^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\tau_{\mathrm{DS}}}}\langle{f}_{{\theta}}(x_{i}),{f}_{{\theta}}(t_{c_{i}})\rangle}}{\sum_{j=1}^{{B^{\prime}}}e^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\tau_{\mathrm{DS}}}}\langle{f}_{{\theta}}(x_{i}),{f}_{{\theta}}(t_{c_{j}})\rangle}}\right)+\sum_{i=1}^{{B^{\prime}}}-\log\left(\frac{e^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\tau_{\mathrm{DS}}}}\langle{f}_{{\theta}}(t_{c_{i}}),{f}_{{\theta}}(x_{i})\rangle}}{\sum_{j=1}^{{B^{\prime}}}e^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\tau_{\mathrm{DS}}}}\langle{f}_{{\theta}}(t_{c_{i}}),{f}_{{\theta}}(x_{j})\rangle}}\right).$
C.2 Experiment configurations
Table 6: Detailed hyper parameters of the different configurations used for training and ablations of FLYT and M-FLYT. * Initialization “DataComp medium” means we train a model from scratch on the unfiltered DataComp medium scale dataset using the DataComp medium scale training configuration.
Parameter FLYT-Base FLYT-22 FLYT-LR M-FLYT-Base
Optimizer AdamW AdamW AdamW AdamW
Weight decay 0.2 0.2 0.2 0.2
$(\beta_{1},\beta_{2})$ $(0.9,0.98)$ $(0.9,0.98)$ $(0.9,0.98)$ $(0.9,0.98)$
Scheduler cosine cosine cosine cosine
Reference model LR 5e-5 5e-5 Variable 5e-5
Scoring model LR 1e-3 1e-3 Variable 1e-3
Scoring model base init OpenAI OpenAI OpenAI None
Reference model init OpenAI OpenAI OpenAI DataComp medium
Downstream loss CE + temperature CLIP CE CE
Downstream batch size ${B^{\prime}}$ 3072 768 1536 3072
Upstream batch size ${B}$ 4096 4096 4096 4096
Num. downstream samples 15.4M 3.8M 7.7M 15.4M
Num. upstream samples 20.5M 20.5M 20.5M 20.5M
Num. steps 5000 5000 5000 5000
Num. warmup steps 100 100 100 100
Downstream dataset ImageNet 22 tasks ImageNet ImageNet
Upstream dataset DataComp medium DataComp medium DataComp medium DataComp medium
Trainable reference parameters all all all all
Trainable scoring parameters only head only head only head all
When training FLYT and M-FLYT, we optimize both the scoring model, the mixing model and the reference models using AdamW [24] with a weight decay of 0.2, $(\beta_{1},\beta_{2})=(0.9,0.98)$, and a cosine scheduler [23]. Due to computational constraints, some experiments compare against model configurations that differ from those listed in Section 4. We list their full hyper parameters in Table 6 and describe their main differences here. We refer to the configurations from Section 4.2, and Section 4.3 as FLYT-Base and M-FLYT-Base respectively. FLYT-22 maintains the same parameters as FLYT-Base, with three key differences: it employs the standard CLIP loss as the downstream loss, uses a downstream batch size ${B^{\prime}}$ of 768, and trains with 22 downstream tasks. FLYT-LR uses the same configuration as FLYT-Base, except it employs the CE loss without temperature and uses a downstream batch size ${B^{\prime}}$ of 1536; this configuration was used exclusively for learning rate experiments.
C.3 FLYT Learning Rate Tuning
Table 7 compares different scoring model and reference model learning rates using the FLYT-LR configuration. We notice that FLYT is not very sensitive to the scoring model LR (with reference LR fixed at 1e-4) and select 1e-3 based on slightly better results. FLYT is somewhat sensitive to reference model LR, and with scoring model LR fixed at 1e-3, 5e-5 gives the best results. These values are used for subsequent experiments.
Table 7: Learning rate tuning using the FLYT-LR configuration (Table 6). We first optimize the scoring model LR with reference model LR fixed at 1e-4, then optimize reference model LR with scoring LR fixed at 1e-3.
Scoring model LR Reference model LR ImageNet Average
5e-4 1e-4 0.286 0.300
1e-3 1e-4 0.286 0.306
5e-3 1e-4 0.280 0.310
1e-2 1e-4 0.277 0.300
1e-3 1e-5 0.158 0.238
1e-3 5e-5 0.292 0.319
1e-3 1e-4 0.286 0.306
1e-3 5e-4 0.267 0.296
C.4 FLYT Ablations
Table 8 presents experiments using FLYT-Base. It shows a comparison between downstream losses, where using CE with a temperature parameter is comparable to the CLIP downstream loss, both outperforming the standard CE loss. It compares reference model and scoring model base initializations, where initializing with stronger models outperform weaker ones. We evaluate four initialization methods in ascending order of performance: Random initialization, “DataComp medium”, where we train a model from scratch on the unfiltered DataComp medium scale dataset using the DataComp medium scale training configuration, OpenAI checkpoint, and “DataComp-13B”, which is a model trained by Gadre et al. [15] on the DC-1B dataset for 13B samples seen. Table 8 also shows that training the reference model is much better than freezing it, but freezing the scoring model base is required for the training to succeed. Our ablation studies show that performance remains stable with downstream batch sizes ranging from 1024 to 4096. Similarly, performance is consistent when training for 1250 to 12500 steps, suggesting that FLYT could be optimized using less compute.
Table 8: Experiments using the FLYT-Base configuration (Table 6). * Initialization “Trained on DC-1B” means using the ViT-B-32 model trained by Gadre et al. [15] on DC-1B for 13B samples seen (Hugging Face model card).
Experiment Baseline Variation ImageNet Average
FLYT-Base 0.316 0.323
Downstream loss CE+temperature CE (no temperature) 0.287 0.305
CLIP 0.313 0.330
Reference model initialization OpenAI Random initialization 0.270 0.312
Trained on DC medium 0.290 0.308
Trained on DC-1B 0.314 0.320
Scoring model base initialization OpenAI Random initialization 0.157 0.234
Trained on DC medium 0.287 0.302
Trained on DC-1B 0.325 0.329
Reference and scoring model base initialization OpenAI Trained on DC-1B 0.326 0.331
Trainable parameters Head + reference model Only head 0.251 0.298
Everything 0.135 0.227
Num. training steps 5000 12500 0.306 0.320
1250 0.311 0.329
Downstream Batch Size 3072 1024 0.315 0.313
4096 0.311 0.316
C.5 FLYT with Diverse Downstream Data
Instead of using the ImageNet training set alone for the downstream data, we experiment with using 22 training sets from the DataComp benchmark. We now detail the training configuration, and then discuss the results of experiments using FLYT-22. First, we list the datasets used:
  • ImageNet [11]
  • MNIST [10]
  • SVHN [29]
  • CIFAR10 [22]
  • CIFAR100 [22]
  • Food-101 [3]
  • Oxford Flowers-102 [30]
  • Oxford-IIIT Pet [31]
  • iWildCam [2]
  • SUN-397 [47]
  • EuroSAT [18]
  • Stanford Cars [21]
  • DTD [8]
  • RESISC45 [5]
  • GTSRB [43]
  • FGVC Aircraft [28]
  • Pascal VOC 2007 [13]
  • Country211 [33, 45]
  • Rendered SST2 [50]
  • FMoW [7]
  • Dollar Street [35]
  • STL-10 [9]
When preparing our datasets, we split them into tar files of 100 examples each in webdataset format. During training, we use a shuffle buffer of 50,000 examples. Since the webdataset loader reads entire tar files at once, small shards and a large shuffle buffer help ensure diverse tasks within each training batch. The combined dataset contains 2M examples, with ImageNet making up the majority. In Table 9, we compare different dataset configurations. Using the combined dataset without balancing classes yields performance very similar to using only the ImageNet dataset. When we balance the datasets to ensure equal representation (where the model sees samples from each downstream dataset the same number of times), performance deteriorates. This performance drop might be due to some tasks containing as few as 1020 total training samples, potentially causing overrepresentation of these smaller datasets during training.
Table 9: Different downstream tasks experiments using the FLYT-22 configuration (Table 6).
Downstream tasks variation ImageNet Average
Imbalanced 22 tasks 0.315 0.330
Balanced 22 tasks 0.274 0.311
Only ImageNet 0.313 0.330
C.6 M-FLYT Ablations
Table 10 shows experiments using M-FLYT. Unlike our results with FLYT, the standard CE loss slightly outperforms CE with temperature, and reference model initialized with “DataComp medium” (explained in Section C.4) outperforms OpenAI initialization. We also find that input standardization greatly increases performance.
Table 10: Experiments using the M-FLYT-Base configuration (Table 6).
Experiment Baseline Variation ImageNet Average
M-FLYT 0.359 0.371
Downstream loss CE CE + temperature 0.352 0.366
Reference model initialization Trained on DC medium Random initialization 0.341 0.354
OpenAI 0.342 0.355
Input standardization Yes No 0.323 0.337
Table 11 shows the result of training M-FLYT while excluding groups of related inputs. Removing any input group degrades performance, with several notable results: DFN reproduction scores have the strongest impact on ImageNet, ImageNet distribution shifts, and retrieval scores (degrading performance by 1.8%, 1.4%, and 2.9% respectively). For VTAB tasks, excluding HYPE scores causes the largest drop (1.7%), while unexpectedly, removing the linear classifiers—despite being the weakest standalone filtering methods—leads to the largest decline in average performance (1.6%).
Table 11: Comparison of training M-FLYT with all input methods vs leaving one method out. Numbers in parentheses show performance degradation compared to using all inputs. We highlight values with the largest degradation, indicating which methods provide the most benefit when included.
Inputs ImageNet ImageNet dist. shifts VTAB Retrieval Average
All Inputs 0.359 0.294 0.383 0.310 0.371
No ViT-B/32, ViT-L/14 0.355 (0.004) 0.288 (0.006) 0.367 (0.016) 0.307 (0.003) 0.360 (0.011)
No DFN-Base, DFN-FT 0.341 (0.018) 0.280 (0.014) 0.375 (0.008) 0.281 (0.029) 0.358 (0.013)
No ${-d_{L}},{\epsilon_{i}},{\epsilon_{t}}$ 0.351 (0.008) 0.291 (0.003) 0.366 (0.017) 0.308 (0.002) 0.364 (0.007)
[…]
No IN1K-Classifier, CC2M-Classifier 0.352 (0.007) 0.288 (0.006) 0.368 (0.015) 0.301 (0.009) 0.355 (0.016)
No FLYT 0.352 (0.007) 0.292 (0.002) 0.376 (0.007) 0.309 (0.001) 0.366 (0.005)
C.7 Comparing M-FLYT to Aggregation Methods in Prior Work
Table 12 compares M-FLYT against the original mixing strategies of HYPE and NegCLIPLoss. For HYPE’s components, M-FLYT outperforms the simple summation approach on both ImageNet (+1.3%) and average (+0.6%) metrics. When mixing NegCLIPLoss scores, M-FLYT performs similarly to the original two-stage filtering method, showing marginal improvement on ImageNet (0.3%) but slightly lower average performance (0.6%).
Table 12: Comparing M-FLYT to the original mixing method of HYPE [20] and NegCLIPLoss [46]. HYPE combines scores by summing their three components (${-d_{L}}+{\epsilon_{i}}+{\epsilon_{t}}$), while NegCLIPLoss applies a two-stage filtering process: selecting the top 30% based on NegCLIPLoss scores, then selecting the top 20% according to their NormSim scores from this filtered subset.
Input method Mixing method ImageNet Average
HYPE Original (sum scores) 0.309 0.341
M-FLYT 0.322 0.347
NegCLIPLoss and NormSim Original (two stage) 0.331 0.363
M-FLYT 0.334 0.357
C.8 DataComp Small Scale Results
For the small scale DataComp benchmark, we reuse the mixing model from the medium scale without retraining FLYT or M-FLYT. The only changes were re-tuning the repetition penalty ${\alpha}$ for the smaller dataset and reducing the sampling batch size ${G}$ by a factor of 10 to $10K$, aligning with the dataset size reduction. Figure 4 shows that the smaller scale performs best with a higher repetition penalty, around ${\alpha}=0.5$, compared to the medium scale which performs better with values between 0.15 to 0.25. A higher repetition penalty reduces the number of repetitions per example, which becomes more important for smaller datasets.
Figure 4: Comparing SCS to hard threshold filtering on the small scale DataComp benchmark. SCS was tested with ${\alpha}$ values from 0.15 to 1, and standard threshold with values from 10% to 30%
Appendix D Reproduction comparison
Table 13: Comparison between the results reported in the original papers and our reproduction using their provided UIDs. Some experiments were run using the complete DataComp medium scale dataset (128M) while our reproduction used the 119M examples we were able to download. Results marked with * were performed by Wang et al. [46] on their 110M downloaded dataset; they also report partial results using the full 128M data pool. Our best results are listed below.
Filtering Method Pool Size ImageNet ImageNet dist. shifts VTAB Retrieval Average
110M 0.324 0.274 0.359 0.263 0.352
NegCLIPLoss 119M (Reprod.) 0.333 0.273 0.361 0.251 0.352
128M 0.371 0.298 0.388 0.288 0.373
DFN 119M (Reprod.) 0.367 0.303 0.371 0.280 0.364
116M 0.382 0.303 0.393 0.306 0.379
HYPE+DFN 119M (Reprod.) 0.372 0.304 0.374 0.287 0.368
128M, 110M 0.375 0.309 0.386 0.281 0.386
DFN+NegCLIPLoss 119M (Reprod.) 0.371 0.304 0.390 0.285 0.378
128M, 110M 0.382 0.314 0.385 0.276 0.388
HYPE+DFN+NegCLIPLoss 119M (Reprod.) 0.381 0.310 0.392 0.284 0.378
M-FLYT+SCS (${\alpha}=0.15$) 119M 0.401 0.311 0.391 0.292 0.375
M-FLYT+SCS (${\alpha}=0.25$) 119M 0.395 0.309 0.393 0.310 0.378
Table 13 compares model we train using publicly available DataComp subsets to the numbers reported in the papers that published them. The results are generally similar, but not identical, with accuracy difference of up to 1%. The main driver of the difference appears to be the size of the actual pool used to train the model: we were able to download only 119M of the 128M examples in the DataComp medium scale pool, while the authors of the other papers had access to a more complete subset of the examples.